# Social Bias Probing: Fairness Benchmarking for Language Models

### v3

@@ 1 Introduction
Our Social Bias Probing framework, as outlined in Fig. 1, is specifically designed to enable a nuanced understanding of biases inherent in language models. Accordingly, the input of our approach consists of a set stereotypes and identities. To this end, we generate our probing dataset by combining stereotypes from the Social Bias Inference Corpus (SBIC; Sap et al. 2020) and identities from the lexicon by Czarnowska et al. (2021). In this paper we examine identities belonging to four social categories: gender, religion, disability, and nationality. Secondly, we assess social biases across five state-of-the-art LMs in English. We use perplexity (Jelinek et al., 1977), a measure of language model uncertainty, as a proxy for bias. By analyzing the variation in perplexity when probes feature different identities from diverse social categories, we infer which identities are deemed most likely by a model. This approach facilitates a three-dimensional analysis – by social category, identity, and stereotype—across the evaluated LMs. In summary, the contributions of this work are:
@@
  • We introduce SoFa (Social Fairness), a benchmark for fairness probing addressing limitations of existing datasets, including a variety of different identities and stereotypes. [^1] SoFa will be made available upon paper acceptance. See the data statement in App. A.
@@ 3.2 Identities
While we could have directly used the identities provided in the SBIC dataset, we chose not to due to their unsuitability from frequent repetitions and varying expressions influenced by individual annotators’ styles. To leverage a coherent set of analyzed identities, we deploy the lexicon [^7] The complete list of identities is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. created by Czarnowska et al. (2021). In Tab. 3 in the Appendix, we report samples for each category. We map the SBIC dataset group categories to the identities available in the lexicon (Tab. 5 in Appendix). Specifically, the categories from SBIC are gender, race, culture, disabilities, victim, social, and body. We first define and rename the culture category to include religions and broaden the scope of the race category to encompass nationalities. We then link the categories in the SBIC dataset to those present in the lexicon as follows: gender identities are drawn from the lexicon’s genders and sexual orientations, nationality from race and country categories, religion and disabilities directly from their respective categories. This mapping excludes the broader SBIC categories–victim, social, and body–due to alignment challenges with lexicon entries and difficulties in preserving statement invariance. [^8] This choice is motivated by the fact that the stereotypes under these categories are often specific to a particular identity; for example, they might have referenced body parts belonging to one gender and not another. While we inherit the assignment of an identity to a specific category the underlying resources, we recognize that these framings may simplify the complexity of identities.
@@ 3.5 Fairness Evaluation
Table 1: Results on SoFa and the two previous fairness benchmarks, StereoSet and CrowS-Pairs. The ranking ranges from 1 (LM most biased) to 10 (LM least biased $\downarrow$); for each of the scores, the best value in bold is the lowest $\downarrow$, connoting the least biased model. We note the number of instances in each dataset next to their names.
@@ 4.2 Results
Table 2: SoFa score by category: best ($\downarrow$) value in bold.
@@
Across all models except for BLOOM-3b, religion consistently stands out as the category with the most pronounced disparity, while nationality often shows the lowest value. Given the extensive focus on gender and racial biases in the NLP literature, it’s plausible that recent language models have undergone some degree of fairness mitigation for these particular biases, which may explain why religion now emerges more prominently. Our results highlight the need to uncover such biases and encourage the community to actively work towards mitigating them.
@@
We display, in Fig. 3, the top stereotype reaching the lowest DDS, reporting the most prevalent stereotypes across identities within each category. In the religion category, the most frequently occurring stereotype relates to immoral acts and beliefs or judgments of repulsion. For the gender category, mentions of stereotypical behaviors and sexual violence are consistently echoed across models, while in the nationality category, references span the lack of employment, physical violence (both endured and performed), and crimes. Stereotypes associated with disability encompass judgments related to appearance, physical incapacity, and other detrimental opinions. We observe that the harms that identities experience in real life, such as sexual violence against women (Russo and Pirlott, 2006; Tavara, 2006), high unemployment of immigrants (discussed in terms of nationalities) (Appel et al., 2015; Olier and Spadavecchia, 2022), and stigmatized appearance of people with disabilities (Harris, 2019), are indeed reflected by the models’ behavior.
@@ 5 Conclusion
Our benchmark is highly reliable on the set of stereotypes and identities included in the probing dataset. We opted to use the list of identities from Czarnowska et al. (2021). However, the identities included encompass a range of perspectives that the lexicon in use may not fully capture. Moreover, the stereotypes we adopt are derived from SBIC, which aggregated potentially biased content from a variety of online platforms such as Reddit, Twitter, and specific hate sites Sap et al. (2020). These platforms tend to be frequented by certain demographics. Despite having a broader demographic than traditional media sources such as newsrooms, Wikipedia editors, or book authors (Wagner et al., 2015), they predominantly reflect the biases and perspectives of white men from Western societies.
@@
In light of these considerations, following Attanasio et al. (2022), we advocate for the responsible use of benchmarking suites. Our benchmark is intended to be a starting point, and we recommend its application in conjunction with human-led evaluations. Users are encouraged to further develop and refine our probing dataset to enhance its inclusivity in terms of identities, stereotypes, and models included.
@@ C.1 Dataset Statistics
In Tab. 3, we report example identities for each category of the SoFa dataset. We deploy the lexicon created by Czarnowska et al. (2021): the complete list is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. Tab. 4 shows a sample of the probes included in our SoFa dataset. In Tab. 5, we document the coverage statistics regarding targeted categories and identities of SoFa. We also include the descriptions of SBIC, StereoSet, and CrowS-Pairs for comparison. Since the categories in SoFa differ and do not correspond to the two competitor datasets, i.e., a one-to-one mapping is absent, we report only quantities for overlapping categories, as we shall specify (for completeness, we indicate in parentheses the full size of their datasets in the total column). To calculate the probes for CrowS-Pairs, we combine the categories of nationality and race/color for Nationality, and the categories of gender/gender identity and sexual orientation for Gender. Lastly, considering that CrowS-Pairs do not encode identities but only categories, we do not include the number of identities per category for this dataset.
Finally, we also report in Tab. 4 the dataset structure along with sample instances from SoFa.
@@ C.2 Stereotype Clustering
We provide an overview of the main stereotype clusters included in SoFa. First, we use gte-base-en-v1.5, a state-of-the-art pre-trained sentence transformer (Li et al., 2023), to produce an embedding for each stereotype. Second, we reduce dimensionality to d = 15 with UMAP (McInnes et al., 2018), to reduce complexity prior to clustering. Third, we cluster the stereotypes using HDBScan (McInnes et al., 2017), a density-based clustering algorithm, which does not force cluster assignment: 57% of prompts are assigned to 15 clusters and 43% are various stereotypes. We use a minimum cluster size of 90, ($\approx$ 1% of 9,102 stereotypes) and a minimum UMAP distance of 0. Other hyperparameters are default. To interpret the identified clusters, we use TF-IDF to extract the top 10 most salient uni- and bigrams from each cluster’s prompts, and locate five prompts closest and furthest to the cluster centroids. Finally, we use GPT-4 to assign a short descriptive name to each cluster based on the top n-grams and closest stereotypes. In Fig. 5, we present a distribution of stereotypes in these clusters.
@@ C.3 Hate Speech Analysis
As expected, the stereotypes of SoFa do not display evident features of Hate Speech since they stand for different, more complex, and nuanced phenomena. Furthermore, we highlight that we do not have a ground truth concerning hatefulness for these stereotypes. Therefore, we must also consider a certain margin of error caused by the classifier in ambiguous or uncertain instances. A more suitable lens for analyzing the contents of this dataset could be harmfulness or hurtfulness Nozza et al. (2021), featured by apparently neutral statements. Harmfulness can be implicit, and it is present in our implied statements, which, as outlined in Section A), express harmful stereotypical beliefs. However, the harmfulness evaluation is more challenging to grasp and still poorly explored. Crucially, stereotypes and Hate Speech are two different phenomena and, as such, need to be investigated and addressed separately, requiring targeted approaches. Indeed, identifying when a stereotype is expressed non-offensively remains a challenge and an ongoing research area Havens et al. (2022).
@@ Appendix D Experimental Setup
In Tab. 6, we list the selected LMs: for each, we examine two scales with respect to the number of parameters.
@@
Appendix E Supplementary Material
Fig. 8 shows correlation heat map between $PPL^{\star}$ of the various LMs and stereotype length. The correlation is negative (below 00) but not extremely high, indicating a weak relationship. Specifically, this means that shorter lengths correspond to higher $PPL^{\star}$. Overall, we recall that the range of lengths is moderate, i.e., reaching a maximum of $14$ words.

### v4

@@ 1 Introduction
Our Social Bias Probing framework, as outlined in Fig. 1, is specifically designed to enable a nuanced understanding of biases inherent in language models. Accordingly, the input of our approach consists of a set stereotypes and identities. To this end, we generate our probing dataset by combining stereotypes from the Social Bias Inference Corpus (SBIC; Sap et al. 2020) and identities from the lexicon by Czarnowska et al. (2021). In this paper we examine identities belonging to four social categories: gender, religion, disability, and nationality. Secondly, we assess social biases across five state-of-the-art LMs in English. We use perplexity (Jelinek et al., 1977), a measure of language model uncertainty, as a proxy for bias. By analyzing the variation in perplexity when probes feature different identities within the diverse social categories, we infer which identities are deemed most likely by a model. This approach facilitates a three-dimensional analysis – by social category, identity, and stereotype—across the evaluated LMs. In summary, the contributions of this work are:
@@
  • We introduce SoFa (Social Fairness), a benchmark for fairness probing addressing limitations of existing datasets, including a variety of different identities and stereotypes. [^1] SoFa is available at https://huggingface.co/datasets/copenlu/sofa. See the Data Statement in App. A.
@@ 3.2 Identities
Although we could have directly used the identities provided in the SBIC dataset, we opted not to, as they were unsuitable due to belonging to multiple overlapping categories and often being repeated in various wording, influenced by the differing styles of individual annotators. To leverage a coherent distinct set of identities, we deploy the lexicon [^7] The complete list of identities is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. created by Czarnowska et al. (2021). In Tab. 3 in the Appendix, we report samples for each category. We map the SBIC dataset group categories to the identities available in the lexicon (Tab. 5 in Appendix). Specifically, the categories from SBIC are gender, race, culture, disabilities, victim, social, and body. We first define and rename the culture category to include religions and broaden the scope of the race category to encompass nationalities. We then link the categories in the SBIC dataset to those present in the lexicon as follows: gender identities are drawn from the lexicon’s genders and sexual orientations, nationality from race and country categories, religion and disabilities directly from their respective categories. This mapping excludes the broader SBIC categories–victim, social, and body–due to alignment challenges with lexicon entries and difficulties in preserving statement invariance. [^8] This choice is motivated by the fact that the stereotypes under these categories are often specific to a particular identity; for example, they might have referenced body parts belonging to one gender and not another. While we inherit the assignment of an identity to a specific category the underlying resources, we recognize that these framings may simplify the complexity of identities.
@@ 3.5 Fairness Evaluation
Table 1: Results on SoFa and the two previous fairness benchmarks, StereoSet and CrowS-Pairs. We recall that while SoFa reports an average of variances, the other two benchmarks feature the scores as percentages. The ranking, which allows a more intuitive comparison of the scores, ranges from 1 (LM most biased) to 10 (LM least biased $\downarrow$); for each of the scores, the best value in bold is the lowest $\downarrow$, connoting the least biased model. We note the number of instances in each dataset next to their names.
@@ 4.2 Results
Table 2: SoFa score reporting an average of variances by category: best ($\downarrow$) value in bold.
@@
Across all models except for BLOOM-3b, religion consistently stands out as the category with the most pronounced disparity, while nationality often shows the lowest value. Given the extensive focus on gender and racial biases in the NLP literature, it is plausible that recent language models have undergone some degree of fairness mitigation for these particular biases, which may explain why religion now emerges more prominently. Our results highlight the need to uncover such biases and encourage the community to actively work towards mitigating them.
@@
We display, in Fig. 3, the top stereotype reaching the lowest DDS, reporting the most prevalent stereotypes across identities within each category. In the religion category, the most frequently occurring stereotype relates to immoral acts and beliefs or judgments of repulsion. For the gender category, mentions of stereotypical behaviors and sexual violence are consistently echoed across models, while in the nationality category, references span the lack of employment, physical violence (both endured and performed), and crimes. Stereotypes associated with disability encompass judgments related to appearance, physical incapacity, and other detrimental opinions.
Overall, we observe that the harms that identities experience in real life, such as sexual violence against women (Russo and Pirlott, 2006; Tavara, 2006), high unemployment of immigrants (discussed in terms of nationalities) (Appel et al., 2015; Olier and Spadavecchia, 2022), and stigmatized appearance of people with disabilities (Harris, 2019), are indeed reflected by the models’ behavior.
@@ 5 Conclusion
Our benchmark is highly reliant on the set of stereotypes and identities included in the probing dataset. We opted to use the list of identities from Czarnowska et al. (2021). However, the identities included encompass a range of perspectives that the lexicon in use may not fully capture. Moreover, the stereotypes we adopt are derived from SBIC, which aggregated potentially biased content from a variety of online platforms such as Reddit, Twitter, and specific hate sites Sap et al. (2020). These platforms tend to be frequented by certain demographics. Despite having a broader demographic than traditional media sources such as newsrooms, Wikipedia editors, or book authors (Wagner et al., 2015), they predominantly reflect the biases and perspectives of white men from Western societies.
@@
In light of these considerations, we advocate for the responsible use of benchmarking suites Attanasio et al. (2022). Our benchmark is intended to be a starting point, and we recommend its application in conjunction with human-led evaluations. Users are encouraged to further develop and refine our dataset to enhance its inclusivity in terms of identities, stereotypes, and models included.
Acknowledgements
This research was co-funded by Independent Research Fund Denmark under grant agreement number 9130-00092B, and supported by the Pioneer Centre for AI, DNRF grant number P1. The work has also been supported by the European Community under the Horizon 2020 programme: G.A. 871042 SoBigData++, ERC-2018-ADG G.A. 834756 XAI, G.A. 952215 TAILOR, PRIN 2022 PIANO (Personalized Interventions Against Online Toxicity) project under CUP B53D23013290006, and the NextGenerationEU programme under the funding schemes PNRR-PE-AI scheme (M4C2, investment 1.3, line on AI) FAIR (Future Artificial Intelligence Research). The first author would like to thank Isacco Beretta for the constructive feedback. Finally, we thank the anonymous reviewers for their helpful suggestions.
@@ C.1 Dataset Statistics
In Tab. 3, we report example identities for each category of the SoFa dataset. We deploy the lexicon created by Czarnowska et al. (2021): the complete list is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. Tab. 4 shows a sample of the probes included in our SoFa dataset. In Tab. 5, we document the coverage statistics regarding targeted categories and identities of SoFa. We also include the descriptions of SBIC, StereoSet, and CrowS-Pairs for comparison. Since the categories in SoFa differ and do not correspond to the two competitor datasets, i.e., a one-to-one mapping is absent, we report only quantities for overlapping categories, as we shall specify (for completeness, we indicate in parentheses the full size of their datasets in the total column). To calculate the probes for CrowS-Pairs, we combine the categories of nationality and race/color for Nationality, and the categories of gender/gender identity and sexual orientation for Gender. Lastly, considering that CrowS-Pairs do not encode identities but only categories, we do not include the number of identities per category for this dataset. Finally, we also report in Tab. 4 the dataset structure along with sample instances from SoFa.
@@ C.2 Stereotype Clustering
We provide an overview of the main stereotype clusters included in SoFa. First, we use gte-base-en-v1.5, a state-of-the-art pre-trained sentence transformer (Li et al., 2023), to produce an embedding for each stereotype. Second, we reduce dimensionality to d = $15$ with UMAP (McInnes et al., 2018), to reduce complexity prior to clustering. Third, we cluster the stereotypes using HDBScan (McInnes et al., 2017), a density-based clustering algorithm, which does not force cluster assignment: $57$% of prompts are assigned to $15$ clusters and $43$% are various stereotypes. We use a minimum cluster size of $90$, ($\approx$ $1$% of $9,102$ stereotypes) and a minimum UMAP distance of 00. Other hyperparameters are default.
To interpret the identified clusters, we use TF-IDF to extract the top 10 most salient uni- and bigrams from each cluster’s prompts, and locate 5 prompts closest and furthest to the cluster centroids. Finally, we use GPT-4 to assign a short descriptive name to each cluster based on the top n-grams and closest stereotypes. See the prompt used below.
[…]
In Fig. 5, we present a distribution of stereotypes in these clusters. Stereotypes associated with sexualization and violence are the most prevalent in SoFa, followed by family neglect, while slavery and sports restrictions are the least common.
@@ C.3 Hate Speech Analysis
As expected, the stereotypes of SoFa do not display evident features of Hate Speech since they stand for different, more complex, and nuanced phenomena. Furthermore, we highlight that we do not have a ground truth concerning hatefulness for these stereotypes. Therefore, we must also consider a certain margin of error caused by the classifier in ambiguous or uncertain instances. A more suitable lens for analyzing the contents of this dataset could be harmfulness or hurtfulness Nozza et al. (2021), featured by apparently neutral statements. Harmfulness can be implicit, and it is present in our implied statements, which, as outlined in Appendix A, express harmful stereotypical beliefs. However, the harmfulness evaluation is more challenging to grasp and still poorly explored. Crucially, stereotypes and hate speech are two different phenomena and, as such, need to be investigated and addressed separately, requiring targeted approaches. Indeed, identifying when a stereotype is expressed non-offensively remains a challenge and an ongoing research area Havens et al. (2022).
@@ Appendix D Experimental Setup
In Tab. 6, we list the LMs: for each, we examine two scales w.r.t. the number of parameters.
Appendix E Supplementary Material
Fig. 9 shows correlation heat map between $PPL^{\star}$ of the various LMs and stereotype length. The correlation is negative but not extremely high, indicating a weak relationship. Specifically, this means that shorter lengths correspond to higher $PPL^{\star}$. We recall that the range of lengths is moderate, i.e., reaching a maximum of $14$ words.